If one guy paints a house in 12 hrs and another paints it in 6hrs .how long would it take them to paint it together
Answer by checkley79(3341)

1/12+1/6=1/X
1/12+2/12=1/X
3/12=1/X CROSS MULTIPLY.
3X=12
X=12/3
X=4 HOURS WORKING TOGETHER.